OpenStudy (anonymous):

A rectangle with its base on the x-axis has its upper vertices on the curve of y=3-x^2, so find the maximum perimeter of such a rectangle.

jigglypuff314 (jigglypuff314):

erm... all I got so far is that domain restriction is 3-x^2=0

OpenStudy (anonymous):

@sourwing

jigglypuff314 (jigglypuff314):

ok, I got it :) had to use this to get the right idea tho xD http://www.algebra.com/algebra/homework/Surface-area/Surface-area.faq.question.310349.html so from that we get that the two points on the parabola that would make up the vertices of the rectangle (x, 3-x^2) and (-x, 3-x^2) so L=2x and W=3-x^2

jigglypuff314 (jigglypuff314):

then Perimeter is P = 2L + 2W plug in = 2(2x) + 2(3-x^2) = 4x + 6 - 2x^2 <- so take derivative of that

OpenStudy (anonymous):

one sec 314

OpenStudy (anonymous):

-4x+4?

jigglypuff314 (jigglypuff314):

yup :) then solve for x for -4x + 4 = 0

OpenStudy (anonymous):

1

jigglypuff314 (jigglypuff314):

yep :) then plug x = 1 back into L = 2x W = 3-x^2 then into P = 2L + 2W

OpenStudy (anonymous):

l = 2(1) = 2 w = 3-1^2 = 3 p=2(2) + 2(3) = 4 + 6 = 10

jigglypuff314 (jigglypuff314):

erm w = 3-1^2 = 3 - 1 = ?

OpenStudy (anonymous):

2?

jigglypuff314 (jigglypuff314):

I think so xD

OpenStudy (anonymous):

so 2 is the answer righT?

jigglypuff314 (jigglypuff314):

mmm na, I was just correcting your "w = 3-1^2 = 3"

OpenStudy (anonymous):

alright, so where i am goin' from here?

jigglypuff314 (jigglypuff314):

l = 2(1) = 2 w = 3-1^2 = 2 p=2(2) + 2(2) = 4 + 4 = 8 <- tah-dah! XD

OpenStudy (anonymous):

is it's 8?

OpenStudy (anonymous):

<3

jigglypuff314 (jigglypuff314):

yeah, I think so :)

OpenStudy (anonymous):

yep